OpenStudy (anonymous):

which set of equations would be used to solve this word problem? An airplane flew 3 hours with a 30 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air. d = 3(r − 30) and d = 2(r + 30) d = 3(r + 30) and d = 2(r − 30) d = 5(r + 30) None of these systems can be used to solve this problem.

OpenStudy (anonymous):

Do you know how to solve it? :)

OpenStudy (anonymous):

no. can you help with that?

OpenStudy (anonymous):

Yes, I'm trying to figure it out.

OpenStudy (perl):

Let r = speed of plane in still air. In a headwind your speed relative to ground is reduced , this is like a boat traveling up against the current. distance = rate * time d = (r - 30)* 3 Then on return trip you have a tailwind, so groundspeed is increased. d = (r + 30)*2

OpenStudy (anonymous):

Multiply 3×30×2=180÷3